# CNSocialDepress: A Chinese Social Media Dataset for Depression Risk Detection and Structured Analysis

### v1

@@ 6 Conclusion
We introduced CNSocialDepress, a new dataset for depression risk analysis using Chinese social media texts. Uniquely, it is the first dataset to include both user-level and text-level labels, combining binary depression indicators with fine-grained six-dimensional analysis annotated by mental health professionals. We also developed an automated data generation pipeline to ease manual annotation and support language model fine-tuning for classification and summarization tasks. We believe CNSocialDepress will prove valuable across a range of applications, advancing early detection and intervention strategies in mental health.
7 Limitations
@@
While CNSocialDepress provides a valuable resource for depression risk analyses in Chinese social media, several limitations should be noted. First, the dataset is sourced exclusively from Weibo, potentially introducing platform-specific demographic biases (e.g., underrepresentation of rural or elderly populations).
@@
Second, linguistic diversity, including dialectal variations and metaphorical expressions common in Chinese depressive discourse, remains insufficiently captured and may limit cross-regional generalizability. Third, although manual expert annotation ensures high quality, it constrains scalability, resulting in a relatively limited dataset size.
@@
Additionally, the current structured analysis focus primarily on depressive symptoms, omitting comorbid mental health conditions (e.g., anxiety) and contextual social factors. Future work will expand data collection to multi-platform sources, integrate dynamic lexicon updates for emerging expressions, and adopt hybrid annotation frameworks (e.g., expert-guided crowdsourcing) to enhance coverage and efficiency. Privacy-preserving techniques and longitudinal data tracking will further strengthen ethical and practical utility.
8 Ethic Statement
@@
This dataset is built upon the SWDD Cai et al. (2023) benchmark released in early 2023. The original data comes from user-level text on the Weibo platform and was annotated by a team of psychology experts according to the DSM-5 symptom standards.
@@
During our work on the re-annotation process, we applied a second round of deep anonymization to meet privacy requirements. This involved systematically removing personal information, geographic locations, and potential identifiers from the original text, ensuring compliance with data protection protocols Benton et al. (2017) and GDPR (General Data Protection Regulation [^8] EUR-Lex GDPR Regulation). We also invited experienced researchers specializing in depression-related psychology to provide professional annotations based on their expertise.
This dataset is mainly used for building early-warning models for depression risk. Through multi-dimensional feature analysis, it detects potential risk signals and can be used to automatically generate structured psychological assessment reports, offering technical support for digital mental health services. It is important to note that this study is strictly limited to the realm of mental health assistance, and any results should only serve as references for professional diagnosis or as a digital tool for self-assessment.

### v2

@@ 6.  Conclusion
We introduce CNSocialDepress, a new dataset for depression risk analysis from Chinese social media. To our knowledge, it is among the first to provide both user-level and text-level labels. The dataset combines binary depression indicators with fine-grained six-dimensional analyses annotated by mental health professionals. We also develop an automated data generation pipeline. It reduces manual annotation effort and supports language model fine-tuning for classification and summarization. We expect CNSocialDepress to be useful for a range of research and applications, supporting early risk identification and downstream mental health interventions.
7.  Limitations
@@
While CNSocialDepress provides a valuable resource for depression risk analysis on Chinese social media, several limitations should be noted. First, the dataset is sourced exclusively from Weibo. This may introduce platform-specific demographic biases, such as the underrepresentation of rural or elderly populations.
@@
Second, linguistic diversity is not fully captured. This includes dialectal variation and metaphorical expressions that are common in Chinese depressive discourse. This limitation may reduce generalizability across regions. Third, expert annotation ensures high quality, but it limits scalability. It also results in a relatively small dataset.
@@
In addition, the current structured analyses focus primarily on depressive symptoms. They omit comorbid mental health conditions (e.g., anxiety) and broader social context. Future work will expand data collection to multiple platforms. We also plan to update lexicons to capture emerging expressions. Finally, we will explore hybrid annotation frameworks (e.g., expert-guided crowdsourcing) to improve coverage and efficiency. Privacy-preserving techniques and longitudinal tracking can further strengthen ethical and practical utility.
8.  Ethics Statement
@@
This dataset is built upon the SWDD benchmark Cai et al. (2023), released in early 2023. The original data consist of user-level posts from Weibo. They were annotated by psychology experts based on DSM-5 criteria.
@@
During re-annotation, we applied a second round of de-identification to protect privacy. We remove personal information, geographic locations, and other potential identifiers from the texts. We follow recommended data protection practices Benton et al. (2017) and comply with the GDPR (General Data Protection Regulation). [^7] https://eur-lex.europa.eu/legal-content/EN/TXT/PDF/?uri=CELEX:02016R0679-20160504 We also recruit experienced researchers in depression-related psychology to provide expert annotations.
This dataset is intended for research on depression risk identification and for building assistive models. It can support multi-dimensional analysis of risk-related signals and the generation of structured summaries. It is not designed for clinical diagnosis, triage, or treatment decisions. Any outputs should be used only as references for professionals or as a self-assessment aid.